OpenStudy (anonymous):

Normalization of the Gaussian wavepacket I haven't studied complex analysis and I can't figure out why I have to include \[e^{i\phi}\]

OpenStudy (anonymous):

The wavepacket is defined as \[\Psi(x)=\Psi_0e^{\frac{ -(x-x_0)^2 }{ 4\sigma^2 }}\]

OpenStudy (anonymous):

So, what I do is to square the wavepacket to give: \[|\Psi_0|^2 \int\limits_{-\infty}^{\infty}e^\frac{ -(x-x_0)^2 }{ 2\sigma^2 }dx=1\] So, I set \[y = \frac{ -(x-x_0) }{ \sqrt2\sigma }\] and \[dx = \sqrt2\sigma dy\] And the probability becomes \[\sqrt2\sigma |\Psi_0|^2 \int\limits_{-\infty}^{-\infty}e^{-y^2}dy=1\] And after replacing the Gaussian integral, I get \[|\Psi_0|^2 = \frac {1}{\sqrt{2 \pi} \sigma }\] And \[|\Psi_0| = \frac {1}{\ (2 \pi \sigma^2 )^{1/4}} \]

OpenStudy (anonymous):

But then I don't know why when replacing \[ \Psi_0 \] back in the wavepacket definition, I'm supposed to also put \[ e^{i \phi } \] in the numerator of the normalization constant.

OpenStudy (anonymous):

The e^iθ is the phase factor and defines phase for the wave packet. It should be there before normalization as every wave function has a phase factor. It doesn't describe the evolution of phase for the wave packet but rather phase offset in the beginning. Since it's a function of θ, which in this case is a constant, it doesn't change with the integration you've shown.

OpenStudy (anonymous):

Right, thank you!

OpenStudy (anonymous):

You're welcome.